CAGD II April 19, 2006 G. Farin

Programming Assignment # 4 due: 5-1

Morph your face mesh to that of a fellow student. Develop the morphing algorithm based on a decimated mesh, but apply the morph to the full mesh.

The key part to the morph is finding corresponding landmark points o the two meshes. Manually pick about 10 (nose tip etc), triangulate those points and then shoot rays from the triangles to the decimated face mesh. This will give you enough corresponding landmark points. Then feed these points into a trivariate B-spline morpher.

Hand in documented results, but no code. Interactive display in class is also fine.